1. An archer shoots an arrow from a bow. Assuming the force exerted by the bow on the arrow is a spring force,

in one dimension

(a) write an expression of the force exerted on the arrow.

(b) solve the equation in (a) and determine the velocity with which the arrow leaves the bow.

(c) In three dimension write the expression of the force exerted by the bow on the arrow.


1
Expert's answer
2021-10-03T13:23:20-0400

(a) The force is


"F=kx."

(b)

"ma=kx, \\\space\\
m\frac{v^2}{2x}=kx,\\\space\\
v=x\sqrt{\frac{2k}{m}}."

(c) The force is


"\vec F=k_1x_1\hat i+k_2x_2\hat j+k_3x_3\hat k."